OpenStudy (heart_offire):

i need help A store is having a sale on 200 DVD movies. Some of the movies are being sold for $10 and some of them are being sold for $12. If the store made $2200 after selling all 200 DVDs, how many $12 DVDs were sold? A) 50 B) 75 C) 100 D) 150

OpenStudy (heart_offire):

@retirEEd

OpenStudy (janu16):

okk have any idea?

OpenStudy (janu16):

if you don't no problem

OpenStudy (janu16):

$10x+$12y= $2200 x+y=200 multiple the bottom equation by -10 So you now $10x+$12y= $2200 -10x-10y=-2000 (the x's cancel out) nad you are left with 2y=200 y=100 so 100 $12 DVD's were sold

OpenStudy (heart_offire):

well i had an idea but i didnt finish it because i didnt know how to proceed

OpenStudy (anonymous):

\[\text{Solve}[10 (200-x)+12 x=2200,x] \]\[x -> 100 \]

OpenStudy (janu16):

ok